dani120:

Given the equation 3x + 25 = 7x − 5, which order of operations completely solves for x?

Shadow:

Hello @dani120

Shadow:

\[3x + 25 = 7x - 5\]

Shadow:

To solve for x, we want to isolate x. We'll be using PEMDAS, or the order of operations. Do you know what those are?

dani120:

no

Shadow:

Parentheses Exponent Multiplication Division Addition Subtraction There are no parentheses and exponents, so we'll just be using the last four.

Shadow:

Lets start with addition. \[3x + 25 = 7x - 5\] Let's add 5 to both sides \[3x + 30 = 7x\] Now instead of four terms, we have only three. What do you think we should do next?